solve by using the quadratics formula
15nSQUARED-3=4n

Answer by stanbon(75887)

15nSQUARED-3=4n
---------
15n^2-4n-3 = 0
n = [4+-sqrt(4^2-4*15*-3)]/30
n = [4+-sqrt(180)]/30
n = [4+-6sqrt5]/30
n = (2/15) +(1/5)sqrt5 or n = (2/15)-(1/5)sqrt5
=============
Cheers,
Stan H.